### v1

# SatFusion: A Unified Framework for Enhancing Satellite IoT Images via Multi-Temporal and Multi-Source Data Fusion

Abstract.
With the rapid advancement of the digital society, the proliferation of satellites in the Satellite Internet of Things (Sat-IoT) has led to the continuous accumulation of large-scale multi-temporal and multi-source images across diverse application scenarios. However, existing methods fail to fully exploit the complementary information embedded in both temporal and source dimensions. For example, Multi-Image Super-Resolution (MISR) enhances reconstruction quality by leveraging temporal complementarity across multiple observations, yet the limited fine-grained texture details in input images constrain its performance. Conversely, pansharpening integrates multi-source images by injecting high-frequency spatial information from panchromatic data, but typically relies on pre-interpolated low-resolution inputs and assumes noise-free alignment—making it highly sensitive to noise and misregistration. To address these issues, we propose SatFusion: A Unified Framework for Enhancing Satellite IoT Images via Multi-Temporal and Multi-Source Data Fusion. Specifically, SatFusion first employs a Multi-Temporal Image Fusion (MTIF) module to achieve deep feature alignment with the panchromatic image. Then, a Multi-Source Image Fusion (MSIF) module injects fine-grained texture information from the panchromatic data. Finally, a Fusion Composition module adaptively integrates the complementary advantages of both modalities while dynamically refining spectral consistency, supervised by a weighted combination of multiple loss functions. Extensive experiments on the WorldStrat, WV3, QB, and GF2 datasets demonstrate that SatFusion significantly improves fusion quality, robustness under challenging conditions, and generalizability to real-world Sat-IoT scenarios. The code is available at: https://github.com/dllgyufei/SatFusion.git.
Satellite IoT, Data Fusion, Multi-Image Super-Resolution, Pansharpening, Image reconstruction
Guanjie Cheng is the corresponding author.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer visionccs: Computer systems organization Sensor networks
1. Introduction
With the rapid development of communication and computing technologies, the Satellite Internet of Things (Sat-IoT) and satellite computing are becoming key components of future communication and intelligent data processing infrastructures (Wang and Li, 2023; Lofqvist and Cano, 2020; Kothari et al., 2020; Kim et al., 2021). By deploying globally distributed satellite constellations, Sat-IoT enables Earth observation and data transmission capabilities on a planetary scale. To further enhance the role of remote sensing imagery in various applications such as ground observation and smart cities (Wang and Li, 2023; Zhang et al., 2022; Leyva-Mayorga et al., 2023), the demand for high-quality satellite images continues to grow. However, due to the constraints on satellite size and weight, the onboard imaging equipment often suffers from limited spatial resolution (Denby and Lucia, 2020). In addition, the inherent physical limitations of imaging sensors prevent them from simultaneously producing images with both high spatial resolution and high spectral resolution (i.e., with more spectral bands) (Zhang et al., 2022), which restricts the further utilization of Sat-IoT imagery.
Low-Earth orbit (LEO) small satellites serve as the key foundation for constructing Sat-IoT systems, offering advantages such as short revisit cycles and low manufacturing costs (Denby and Lucia, 2020). These capabilities enable multiple satellites to frequently observe the same region, thereby acquiring a large volume of multi-temporal remote sensing images. Meanwhile, the use of different onboard imaging sensors across satellites leads to the generation of multi-source remote sensing images. As the number of observation satellites in Sat-IoT systems continues to grow, ground stations, as receiving terminals, must simultaneously process massive amounts of multi-temporal and multi-source imagery (Yang et al., 2024; Wang et al., 2021; Herrería-Alonso et al., 2023; Wang and Li, 2023; Gao et al., 2020; Li et al., 2022). The contradiction between the abundance of low-quality redundant data and the growing demand for high-quality imagery has become increasingly pronounced (Zhang et al., 2022; Li and Xu, 2023), resulting in severe data backlog issues. Therefore, how to generate high-quality images from massive low-quality and redundant data has emerged as an urgent challenge to be addressed.
Multi-temporal data contain temporal complementary information (Deudon et al., 2020), while multi-source images provide highly complementary spatial and spectral information (Gao et al., 2020). Previous studies have conducted extensive research on multi-temporal data fusion (Multi-Image Super-Resolution, MISR) and multi-source data fusion (Pansharpening). However, existing approaches fail to simultaneously leverage the complementary information from both multi-temporal and multi-source images, leaving several key challenges unresolved: (1) Task separation and lack of a unified framework: The MISR and pansharpening networks are modeled independently (Gao et al., 2020; Li et al., 2022), making it difficult to jointly process multi-temporal and multi-source images in Sat-IoT scenarios. (2) Limited improvement in fusion and enhancement quality: In MISR, the input images usually contain limited fine-grained texture features, which constrains the quality of reconstructed outputs. (3) Significant gap between experimental settings and real satellite applications: Existing pansharpening methods typically construct synthetic datasets using the Wald protocol (Wald et al., 1997), where a single low-resolution multispectral image and a single high-resolution panchromatic image are spatially well-aligned and free from common noise in real-world conditions. Moreover, pansharpening relies on interpolating low-resolution images and spatially aligning them with panchromatic images, which makes it highly sensitive to input quality and thus difficult to apply effectively in practice.
To address the aforementioned background and challenges, we propose SatFusion, a novel unified framework that collaboratively utilizes multi-temporal and multi-source images to reconstruct high-quality imagery. The framework employs a Multi-Temporal Image Fusion (MTIF) module to generate high–spatial-resolution deep features containing temporal complementary information, thereby achieving spatial alignment with the panchromatic image. By leveraging multi-temporal data, this module alleviates the limitations of pansharpening methods that rely on interpolation-based spatial alignment, which are often sensitive to input variations and difficult to apply in real-world conditions, thus enhancing robustness. The Multi-Source Image Fusion (MSIF) module injects fine-grained texture features from the panchromatic image to address the constrained output quality of MISR networks. Furthermore, a Fusion Composition Module adaptively integrates the complementary advantages of multi-temporal and multi-source images while adjusting spectral features, with reconstruction supervised by a weighted combination of multiple loss functions. The main technical contributions of this work are as follows:
(1) We propose a novel unified framework, SatFusion, which collaboratively utilizes multi-temporal and multi-source images to reconstruct high-quality imagery. To the best of our knowledge, this is the first work in the remote sensing domain that jointly optimizes multi-temporal and multi-source image fusion within a unified framework.
(2) As illustrated in Figure  1, SatFusion effectively bridges the gap between massive low-quality redundant data and the growing demand for high-quality imagery, providing an efficient and practical solution for Sat-IoT scenarios that require high-quality imaging.
(3) Extensive experiments on datasets including WorldStrat, WV3, QB, and GF2 demonstrate the superior reconstruction quality and robustness of SatFusion compared with existing methods.
Figure 1. Overview of the SatFusion system. LEO satellites transmit low-quality images to ground stations, where SatFusion performs fusion-based reconstruction to enhance image quality and reduce redundancy. :
2. Background
2.1. Multi-Temporal Data Fusion
Multi-temporal data fusion aims to integrate complementary information from multiple images of the same region captured at different times by the same sensor, thereby reconstructing a higher-quality image. A representative approach is Multi-Image Super-Resolution (MISR). As shown in Figure 2(a), the reconstruction process is formulated as:
(1) $SR=f_{\text{MISR}}(lr_{1},lr_{2},\dots,lr_{T}),$
where the input set $\{lr_{t}\}_{t=1}^{T}\in\mathbb{R}^{\frac{H}{\gamma}\times\frac{W}{\gamma}\times C}$ consists of $T$ low-resolution (LR) images, and the output $SR\in\mathbb{R}^{H\times W\times C}$ denotes a single high-resolution (HR) image. Here, $H$ and $W$ are spatial dimensions, $\gamma$ is the upsampling factor, and $C$ is the number of spectral channels.
In recent years, deep neural network–based methods have demonstrated outstanding performance in terms of accuracy and representational capacity. These include feature alignment based on convolutional neural networks (CNNs) (Deudon et al., 2020; Zhang et al., 2018; Razzak et al., 2023; Salvetti et al., 2020; Valsesia and Magli, 2021), detail enhancement using generative adversarial networks (GANs) (Wolters et al., 2023), and cross-temporal modeling with Transformers (An et al., 2022; Wu et al., 2024; Vaswani et al., 2017; Dosovitskiy et al., 2020; Liu et al., 2021). Nevertheless, since all LR inputs originate from the same modality, they lack sufficient high-frequency texture cues, which fundamentally limits the reconstruction fidelity.
$2.$ Multi-Source Data Fusion
Multi-source data fusion aims to integrate complementary information from images of the same region captured by heterogeneous sensors, thereby reconstructing a higher-quality image. A representative task in this field is pansharpening (Sharpen). As shown in Figure 2(b), the fusion process can be formulated as:
(2) $SR=f_{\text{Sharpen}}(lrms,Pan),$
where $lrms\in\mathbb{R}^{\frac{H}{\gamma}\times\frac{W}{\gamma}\times(C_{ms}>1)}$ denotes a low-resolution multispectral (LRMS) image, and $Pan\in\mathbb{R}^{{H}\times{W}\times(C_{pan}=1)}$ denotes a high-resolution panchromatic (HRPAN) image. The output $SR\in\mathbb{R}^{{H}\times{W}\times C_{ms}}$ is a high-resolution multispectral (HRMS) image with the same spectral channels as the MS input but upsampled by a factor of $\gamma$ in spatial resolution. The introduction of CNNs (Zheng et al., 2020; Duan et al., 2024; Ciotola et al., 2022; Jin et al., 2022), GANs (Zhu et al., 2023), Transformers (Meng et al., 2022; Zhou et al., 2022; Bandara and Patel, 2022), diffusion models (Meng et al., 2023; Xing et al., 2024), and more recently Mamba architectures (He et al., 2025) has significantly improved the fusion accuracy and fidelity of pansharpening networks.
Despite these advances, existing pansharpening networks still face key limitations in real-world Sat-IoT settings. First, most methods rely on a single LRMS and single HRPAN input pair, lacking the ability to exploit abundant multi-temporal observations. Second, since the pioneering PNN model (Masi et al., 2016), nearly all subsequent works have followed the Wald protocol for dataset synthesis, which assumes perfect spatial alignment and noise-free conditions (Wald et al., 1997). This setting diverges from practical Sat-IoT scenarios, where misalignment, blur, and sensor noise are common. Finally, as shown in Figure  2(b), most pansharpening networks first interpolate the input lrmslrms image to match the spatial dimensions of the PanPan image before passing them to the fusion module. The fused features are then processed through either residual connections (Figure  2(d)) or direct outputs (Figure  2(c)) to produce SRSR. Such dependence on a single interpolated input makes these models highly sensitive to LRMS image quality, further limiting robustness under real-world acquisition variability.
Figure 2. Overview of MISR and Pansharpening network general methods. (a) General methods of MISR networks; (b) General methods of Pansharpening networks; (c) Common architecture I of Pansharpening fusion module; (d) Common architecture II of Pansharpening fusion module. :
2.3. Data Processing Methods in Sat-IoT
In the Sat-IoT domain, data selection and data fusion are two primary strategies for managing the massive volume of remote sensing data. Data selection aims to identify high-quality samples, mitigating the conflict between redundant low-quality data and the demand for high-quality imagery. Early methods relied on manually defined rules, such as clarity or cloud coverage thresholds (Zhu and Woodcock, 2012). More recent studies leverage deep learning–based image quality assessment models (Hoeser and Kuenzer, 2020; Hoeser et al., 2020; Jeppesen et al., 2019; Maskey and Cho, 2020), where CNNs automatically learn discriminative features correlated with visual quality, achieving more accurate filtering. However, these approaches are limited to redundancy elimination and cannot synthesize new high-quality data, which constrains their contribution to overall data enhancement.
As discussed in Sections 2.1 and 2.2, data fusion aims to generate higher-quality images from existing datasets. With the rapid growth of Sat-IoT and the accumulation of both multi-temporal and multi-source imagery, exploiting the complementary relationships among heterogeneous datasets has become increasingly promising for real-world applications. To the best of our knowledge, only one concurrent work, (Ziaja et al., 2025), attempts to jointly utilize multi-temporal and multi-source imagery. However, their approach differs fundamentally from ours in both architectural design and motivation. Specifically, Ziaja et al. (2025) trained separate MISR and pansharpening networks and concatenated their outputs sequentially, without addressing the architectural limitations inherent to either model. This two-stage design lacks mutual feature interaction and fails to achieve unified optimization. In addition, their method focuses primarily on compounded resolution enhancement—where the MISR output ($\uparrow\times\alpha$) and HRPAN image serve as inputs to a pansharpening model that still relies on interpolation of low-resolution images for spatial alignment and generates a single super-resolved output ($\uparrow\times\beta$), yielding an overall gain of ($\uparrow\times\alpha\cdot\beta$)—rather than addressing the robustness, noise sensitivity, and scalability challenges characteristic of practical Sat-IoT environments.
Figure 3. Architecture of SatFusion. The framework takes multiple multi-temporal LRMS images and a single HRPAN image as inputs. These are processed sequentially through the multi-temporal image fusion module, the multi-source image fusion module, and the fusion composition module, producing the final HRMS image. :
3. Problem Statement
3.1. Multi-Temporal and Multi-Source Image Fusion Framework
Previous studies cannot simultaneously exploit both multi-temporal and multi-source images in Sat-IoT. We propose a new problem that better fits real-world needs: how to leverage multi-temporal data (a set of LRMS images of the same location captured at different times or by different MS sensors) together with multi-source data (a single HRPAN image and LRMS images of the same location captured by different sensors) to generate a single HRMS image of high quality. The image reconstruction task is formally defined as:
(3) $SR=f_{\text{Reconstruct}}(lrms_{1},lrms_{2},\dots,lrms_{T},Pan),$
where the inputs $\{lrms_{t}\}_{t=1}^{T}\in\mathbb{R}^{\frac{H}{\gamma}\times\frac{W}{\gamma}\times C_{ms}}$ denote LRMS images, and $Pan\in\mathbb{R}^{{H}\times{W}\times C_{pan}}$ denotes one HRPAN image. The output is $SR\in\mathbb{R}^{{H}\times{W}\times C_{ms}}$, representing a HRMS image. The output image preserves the same number of spectral channels as the MS inputs, while being upsampled by a factor of $\gamma$ in both spatial dimensions.
3.2. Design Challenges
Based on the problem formulation above, we identify two major challenges in designing a fusion framework. The first challenge lies in the fact that existing frameworks cannot jointly optimize multi-temporal and multi-source data. As discussed in Section 2, current frameworks each have their own limitations. Constructing a unified framework that organically integrates the optimization of multi-temporal data and multi-source data is nontrivial. Simply training existing multi-temporal and multi-source fusion networks separately and concatenating their outputs in sequence (Ziaja et al., 2025), or entirely discarding prior work, is insufficient.
The second challenge is the need for greater robustness. In real Sat-IoT systems, images often suffer from blur, noise, and imperfect alignment across acquisitions. Most existing studies assume clean and well-aligned inputs, focusing on architectural improvements to optimize performance metrics, which limits their applicability in practice. Our goal is to design a model that can maintain stable performance even under complex and imperfect conditions.
4. Methodology
4.1. Overall Architecture of SatFusion
SatFusion is designed to achieve the following objectives: (1) Construct a unified framework to jointly process multi-temporal and multi-source imagery; (2) Exploit the complementary information of multi-temporal and multi-source images to enhance the quality of the fused output; (3) Maintain robustness under complex real-world conditions; (4) Ensure scalability, such that the framework can be adapted to and extended from prior studies.
SatFusion takes as input a set of LRMS images $\{lrms_{t}\}_{t=1}^{T}\in\mathbb{R}^{\frac{H}{\gamma}\times\frac{W}{\gamma}\times C_{ms}}$ and a single HRPAN image $Pan\in\mathbb{R}^{{H}\times{W}\times(C_{pan}=1)}$. The output is a HRMS fused image $SR\in\mathbb{R}^{{H}\times{W}\times C_{ms}}$ that preserves both high spatial and spectral resolution.
The SatFusion framework consists of three key submodules: a multi-temporal image fusion module, a multi-source image fusion module, and a fusion composition module. The complete network structure is illustrated in Figure  3.
4.2. Implementation Details of Core Modules
Multi-temporal Image Fusion Module (MTIF). Its goal is to extract complementary information from multiple LRMS images $\{lrms_{t}\}_{t=1}^{T}\in\mathbb{R}^{\frac{H}{\gamma}\times\frac{W}{\gamma}\times C_{ms}}$ of the same region acquired by the same type of sensor, so as to generate a single super-resolved MS feature map $SR_{MTIF}\in\mathbb{R}^{H\times W\times C_{ms}}$, which is formally written as:
(4) $SR_{MTIF}=MTIF(lrms_{1},lrms_{2},\dots,lrms_{T}),$
The MTIF module consists of an encoder, a multi-temporal image super-resolution(MISR) module, and a decoder. $\{lrms_{t}\}_{t=1}^{T}$} are passed through the encoder with shared parameters to extract their respective features $\{x_{t}^{encode}\}_{t=1}^{T}$}:
(5) $x_{t}^{encode}=Encoder(lrms_{t}),$
where the output features $\{X_{t}^{encode}\}_{t=1}^{T}\in\mathbb{R}^{\frac{H}{\gamma}\times\frac{W}{\gamma}\times C_{hidden\_encode}}$}, EncoderEncoder denotes the convolutional encoder with shared parameters, and $T$ represents the number of multi-temporal images. Then, $\{X_{t}^{encode}\}_{t=1}^{T}$ are processed by the MISR module to output the deep features of multi-temporal image fusion:
(6) $x^{MISR}=f_{MSIR}(x_{1}^{encode},x_{2}^{encode},\dots,x_{T}^{encode}),$
where the output fused features $x^{MISR}\in\mathbb{R}^{{H}\times{W}\times C_{MISR}}$}, and $f_{MSIR}$ denotes the fusion module of the MISR network designed by previous researchers (MISR module, as shown in Figure 2(a)). Then $x^{MISR}$ is processed by the decoder, and the output is the super-resolved deep feature $SR_{MTIF}$:
(7) $SR_{MTIF}=Decoder(x^{MISR}),$
where the output super-resolved feature $SR_{MTIF}\in\mathbb{R}^{H\times W\times C_{ms}}$}. The function of the decoder is to adjust the spatial feature dimensions of the image so that it aligns with the input HRPAN image PanPan; and to adjust the channel feature dimension of the image so that it matches the target image. We implement the decoder using a sub-pixel convolution (PixelShuffle) block:
(8) $x^{PixelShuffleBlock}=PixelShuffleBlock(Conv2d(x^{MISR})),$
where $x^{PixelShuffleBlock}\in\mathbb{R}^{H^{{}^{\prime}}\times W^{{}^{\prime}}\times C_{ms}}$ is the feature map output by the PixelShuffle Block, and $Conv2d$ adjusts the channel dimension to be divisible by ${\gamma}^{2}$, ensuring the correct operation of the sub-pixel convolution module. The sub-pixel convolution increases the spatial resolution of the feature maps by rearranging pixels from the channel dimension into the spatial dimension. When data enter the PixelShuffle Block, pixel rearrangement is performed first, followed by a stack of Conv, Batch Normalization, and PReLU functions to obtain the output. $x^{PixelShuffleBlock}$ is optionally processed by the ResizeResize module:
(9) $SR_{MTIF}=\begin{cases}x^{PixelShuffleBlock},&\text{if }H=H^{{}^{\prime}}\text{ and }W=W^{{}^{\prime}}\\ Resize(x^{PixelShuffleBlock}),&\text{otherwise}\end{cases}$
where ResizeResize adjusts the spatial dimensions by interpolation to ensure that the output features of the MTIF module are spatially aligned with the input PanPan.
Multi-Source Image Fusion Module (MSIF). Its goal is to extract complementary information from multi-source images of the same region acquired by different sensors, thereby simultaneously preserving spatial details and spectral information. The inputs are the output $SR_{MTIF}$ from the MTIF module and a single HRPAN image $Pan\in\mathbb{R}^{H\times W\times(C_{pan}=1)}$. The output is another super-resolved MS feature map $SR_{MSIF}$, formulated as:
(10) $SR_{MSIF}=MSIF(SR_{MTIF},Pan),$
The MSIF module is composed of a pansharpening module. By feeding $SR_{MTIF}$ and PanPan into the pansharpening module, the fused features $SR_{MSIF}$ are obtained:
(11) $SR_{MSIF}=f_{Sharpen}(SR_{MTIF},Pan),$
Here, the fused features $SR_{MSIF}\in\mathbb{R}^{H\times W\times C_{ms}}$, and $f_{Sharpen}$ represents the fusion module of the pansharpening network designed by previous researchers (Sharpen module, as shown in Figure 2(c)(d)). The key difference from prior work lies in the fact that we take the multi-temporal fused features $SR_{MSIF}$ together with PanPan as inputs to the sharpen module, where the spatial dimensions of $SR_{MSIF}$ and PanPan are already aligned, rather than merely upsampling a single lrmslrms band by interpolation to match the resolution of PanPan.
Fusion Composition Module. Its goal is to fuse the outputs of the MTIF module and the MSIF module, so as to adaptively exploit the advantages of both multi-temporal and multi-source image fusion. The inputs are the output $SR_{MTIF}$ from the MTIF module and the output $SR_{MSIF}$ from the MSIF module, and the output is the final single HRMS image SRSR. The Fusion composition module consists of residual connections and convolution adjustment operations, and finally outputs SRSR:
(12) $SR=Conv(ReLU(Conv(SR_{MTIF}+SR_{MSIF}))).$
Here, the output is the HRMS image $SR\in\mathbb{R}^{H\times W\times C_{ms}}$, and ConvConv denotes a 2D convolution with kernel size 111*1. The 111*1 convolution plays a role in adjusting spectral characteristics (Razzak et al., 2023). By placing the convolution operation after the residual connection, the model is encouraged to learn spectral features more effectively.
4.3. Loss Function
We adopt a supervised learning approach to train SatFusion in an end-to-end manner. We design a weighted combination of different loss functions to supervise the reconstruction results, jointly optimizing both texture features and spectral features of the images:
(13) $\begin{split}\mathcal{L}_{\text{total}}=&\ \lambda_{1}\cdot\mathcal{L}_{\mathrm{MAE}}(\mathit{SR},\mathit{GT})+\lambda_{2}\cdot\mathcal{L}_{\mathrm{MSE}}(\mathit{SR},\mathit{GT})\\ &+\lambda_{3}\cdot\mathcal{L}_{\mathrm{SSIM}}(\mathit{SR},\mathit{GT})+\lambda_{4}\cdot\mathcal{L}_{\mathrm{SAM}}(\mathit{SR},\mathit{GT})\end{split}$
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ are the corresponding weights, satisfying $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}=1$. SRSR denotes the image output by SatFusion, and GTGT denotes the ground truth image. $\mathcal{L}_{\mathrm{MAE}}$ is the $l_{1}$ loss, and $\mathcal{L}_{\mathrm{MSE}}$ is the $l_{2}$ loss, both of which are commonly used to supervise the outputs of MISR and pansharpening networks. Prior works retained the same loss functions while modifying model architectures to ensure fair comparison. On top of these loss functions, we introduce new supervisory signals: $\mathcal{L}_{\mathrm{SSIM}}$ is a texture loss function focusing on the learning of spatial features, while $\mathcal{L}_{\mathrm{SAM}}$ is a spectral loss function emphasizing the learning of spectral features.
Since discrepancies exist between the acquisition conditions of input images and the ground truth, SRSR and GTGT may exhibit horizontal/vertical shifts as well as overall brightness deviations. Following prior work, we perform brightness adjustment and boundary cropping before computing the loss in some experiments (An et al., 2022; Deudon et al., 2020; Salvetti et al., 2020). For more details on SatFusion’s loss calculation, please refer to Appendix D. It is worth emphasizing that we ensure consistency in loss calculation during experiments. Specifically, when comparing SatFusion with baseline methods, we apply the same adjustment operations to SRSR before computing the loss, and use the same functions for loss computation across the entire training, validation, and testing pipeline. This guarantees that improvements in evaluation metrics directly demonstrate the effectiveness of the architecture.
5. Experiments
We conduct extensive experiments to address the questions:
RQ1: How does SatFusion perform compared with baseline models on datasets collected from real-world scenarios?
RQ2: How does SatFusion perform compared with baseline models on synthetic datasets constructed following the Wald protocol?
RQ3: How do the number of input multi-temporal images and the super-resolution scaling factor influence the performance of SatFusion?
RQ4: How do the loss function and the Fusion Module affect the performance of SatFusion?
RQ5: How does SatFusion perform compared with baseline models under challenging conditions such as image blurriness and large cross-modal discrepancies?
We answer RQ1 and RQ2 in the main experiments (Section 5.2), RQ3 in the hyperparameter experiments (Appendix A), RQ4 in the ablation studies (Appendix B), and RQ5 in the robustness experiments (Appendix C).
Table 1. Metrics on the WorldStrat Real (a) and Synthetic (b) datasets.
Methods (a) Metrics on the WorldStrat Real Dataset (b) Metrics on the WorldStrat Synthetic Dataset #Params
[…]
MISR     SRCNN (Dong et al., 2014) 36.963 7536.963\,75 0.876 110.876\,11 2.406 612.406\,61 9.015 049.015\,04 1.457 901.457\,90 9.666 709.666\,70 38.088 0138.088\,01 0.872 100.872\,10 2.417 322.417\,32 6.395 266.395\,26 $1.1857$ $5.7754$ 1778.77K
    HighRes-Net (Deudon et al., 2020) 37.105 7037.105\,70 0.879 490.879\,49 2.232 812.232\,81 9.442 639.442\,63 1.435 501.435\,50 9.404 109.404\,10 38.295 2338.295\,23 0.877 350.877\,35 2.105 352.105\,35 $6.1646$ $1.1363$ $5.2376$ 1631.44K
    RAMS (Salvetti et al., 2020) 37.047 9037.047\,90 0.877 430.877\,43 2.274 102.274\,10 8.981 268.981\,26 1.441 501.441\,50 9.353 609.353\,60 40.469 8140.469\,81 0.920 140.920\,14 1.689 641.689\,64 4.692 994.692\,99 0.882 240.882\,24 $3.0800$ 492.68K
    TR-MISR (An et al., 2022) 37.076 8937.076\,89 0.878 480.878\,48 2.251 362.251\,36 9.255 369.255\,36 1.420 401.420\,40 9.192 909.192\,90 38.760 2538.760\,25 0.886 470.886\,47 1.845 281.845\,28 5.835 875.835\,87 $1.0752$ $4.7981$ 618.00K
Sharpen     PNN (Masi et al., 2016) 46.060 7346.060\,73 0.986 430.986\,43 2.288 562.288\,56 2.787 072.787\,07 0.448 940.448\,94 0.561 220.561\,22 46.742 8446.742\,84 0.991 700.991\,70 1.881 801.881\,80 2.296 762.296\,76 0.409 760.409\,76 0.450 890.450\,89 230.67K
    PanNet (Yang et al., 2017) 46.150 7346.150\,73 0.984 700.984\,70 2.198 172.198\,17 2.756 012.756\,01 0.444 350.444\,35 0.546 970.546\,97 46.873 5746.873\,57 0.989 690.989\,69 1.917 071.917\,07 2.221 952.221\,95 0.394 560.394\,56 0.416 580.416\,58 463.31K
    INNformer (Zhou et al., 2022) 46.241 9546.241\,95 0.985 250.985\,25 2.274 202.274\,20 2.703 452.703\,45 0.437 420.437\,42 0.521 290.521\,29 46.965 0846.965\,08 0.989 650.989\,65 2.078 562.078\,56 2.259 082.259\,08 0.396 350.396\,35 0.437 900.437\,90 218.92K
    Pan-Mamba (He et al., 2025) 45.107 5445.107\,54 0.984 290.984\,29 2.945 152.945\,15 3.022 253.022\,25 0.486 430.486\,43 0.592 590.592\,59 47.319 8147.319\,81 0.992 290.992\,29 1.890 111.890\,11 2.141 302.141\,30 0.381 720.381\,72 0.399 670.399\,67 634.10K
SatFusion MISR Sharpen
SRCNN PNN 47.231 8547.231\,85 0.989 780.989\,78 2.008 162.008\,16 2.371 852.371\,85 0.387 640.387\,64 0.428 780.428\,78 47.604 4847.604\,48 0.991 640.991\,64 1.832 661.832\,66 2.097 712.097\,71 0.364 830.364\,83 0.363 830.363\,83 1853.20K
PanNet 47.323 5447.323\,54 0.989 990.989\,99 1.932 201.932\,20 2.316 972.316\,97 0.384 260.384\,26 0.424 000.424\,$0$ 48.021 3048.021\,30 0.991 210.991\,21 1.735 921.735\,92 2.056 732.056\,73 0.350 500.350\,50 0.345 130.345\,13 2085.84K
INNformer 47.222 1947.222\,19 0.987 200.987\,20 2.051 032.051\,03 2.320 992.320\,99 0.383 400.383\,40 0.418 650.418\,65 47.604 5547.604\,55 0.988 340.988\,34 1.838 461.838\,46 2.134 752.134\,75 0.364 130.364\,13 0.376 880.376\,88 1841.45K
Pan-Mamba 45.629 7945.629\,79 0.988 910.988\,91 2.503 722.503\,72 2.881 162.881\,16 0.473 800.473\,80 0.582 210.582\,21 46.850 5846.850\,58 0.992 050.992\,05 2.072 842.072\,84 2.369 502.369\,50 0.401 640.401\,64 0.435 150.435\,15 2256.63K
HighRes-Net PNN 47.383 4647.383\,46 0.9914 1.936 841.936\,84 2.330 362.330\,36 0.381 720.381\,72 0.423 130.423\,13 48.074 2748.074\,27 0.9928 1.743 341.743\,34 2.046 592.046\,59 0.347 190.347\,19 0.342 130.342\,13 1705.87K
PanNet 47.391 1047.391\,10 0.990 210.990\,21 1.950 811.950\,81 2.323 232.323\,23 0.381 890.381\,89 0.419 290.419\,29 48.065 6948.065\,69 0.991 050.991\,05 1.709 381.709\,38 2.066 132.066\,13 0.344 320.344\,32 0.331 320.331\,32 1938.51K
INNformer 47.402 1147.402\,11 0.990 170.990\,17 1.895 821.895\,82 2.3094 0.383 430.383\,43 0.420 150.420\,15 47.993 1247.993\,12 0.992 750.992\,75 1.754 641.754\,64 2.011 882.011\,88 0.348 840.348\,84 0.346 630.346\,63 1694.12K
Pan-Mamba 46.311 1546.311\,15 0.988 670.988\,67 2.022 892.022\,89 2.602 562.602\,56 0.433 830.433\,83 0.540 480.540\,48 47.717 1047.717\,10 0.989 490.989\,49 1.843 021.843\,02 2.108 762.108\,76 0.360 970.360\,97 0.366 560.366\,56 2109.30K
RAMS PNN 47.452 4047.452\,40 0.990 260.990\,26 1.916 791.916\,79 2.310 852.310\,85 0.378 860.378\,86 0.415 370.415\,37 48.075 8148.075\,81 0.991 580.991\,58 1.688 281.688\,28 2.000 322.000\,32 0.346 240.346\,24 0.333 320.333\,32 567.12K
PanNet 47.076 3547.076\,35 0.988 420.988\,42 1.957 061.957\,06 2.552 612.552\,61 0.394 100.394\,10 0.444 550.444\,55 48.297 0548.297\,05 0.990 620.990\,62 1.6590 1.959 291.959\,29 0.334 760.334\,76 0.311 540.311\,54 799.76K
INNformer 47.4650 0.990 210.990\,21 1.8877 2.364 372.364\,37 0.3766 0.4077 48.4220 0.986 940.986\,94 1.693 471.693\,47 1.9476 0.3304 0.3066 555.37K
Pan-Mamba 46.708 0546.708\,05 0.985 810.985\,81 2.104 862.104\,86 2.538 382.538\,38 0.413 980.413\,98 0.494 130.494\,13 47.810 3847.810\,38 0.987 370.987\,37 1.789 211.789\,21 2.064 722.064\,72 0.355 270.355\,27 0.345 130.345\,13 970.55K
TR-MISR PNN 47.179 4347.179\,43 0.990 480.990\,48 1.914 911.914\,91 2.322 342.322\,34 0.392 030.392\,03 0.445 010.445\,01 47.750 8447.750\,84 $0.9908$ $1.7032$ 2.027 832.027\,83 0.360 810.360\,81 0.367 050.367\,05 692.43K
PanNet 47.159 3447.159\,34 0.990 090.990\,09 1.900 061.900\,06 2.360 062.360\,06 0.394 400.394\,40 0.442 690.442\,69 47.919 9447.919\,94 0.991 120.991\,12 1.690 871.690\,87 1.986 761.986\,76 0.354 940.354\,94 0.365 070.365\,07 925.08K
INNformer 47.172 8147.172\,81 0.986 290.986\,29 1.897 801.897\,80 2.406 592.406\,59 0.391 320.391\,32 0.434 720.434\,72 48.217 2048.217\,20 0.989 540.989\,54 1.745 171.745\,17 1.989 001.989\,$0$ 0.339 070.339\,07 0.329 690.329\,69 680.69K
Pan-Mamba 47.093 3347.093\,33 0.987 930.987\,93 2.005 482.005\,48 2.487 952.487\,95 0.394 280.394\,28 0.449 070.449\,07 47.577 1147.577\,11 0.991 150.991\,15 1.894 161.894\,16 2.169 832.169\,83 0.367 890.367\,89 0.375 820.375\,82 1095.87K
  • The bold entities represent the best results in each sub-block.
Table 2. Experimental metrics on WV3, GF2, and QB datasets
Methods WV3 GF2 QB
[…]
Sharpen     PNN (Masi et al., 2016) $35.9670$ $0.9504$ $4.6518$ $3.9014$ $40.5694$ $0.9641$ $1.9101$ $1.6791$ $35.7257$ $0.9222$ $5.5111$ $5.9618$
    DiCNN (He et al., 2019) $36.3114$ $0.9541$ $4.4815$ $3.7504$ $40.7252$ $0.9658$ $1.8568$ $1.6470$ $35.6112$ $0.9211$ $5.3824$ $6.0262$
    MSDCNN (Yuan et al., 2018) $35.3417$ $0.9399$ $5.1254$ $4.1817$ $40.5558$ $0.9643$ $1.9400$ $1.6826$ $35.5545$ $0.9201$ $5.4935$ $6.0724$
    DRPNN (Wei et al., 2017) $36.5765$ $0.9562$ $4.4941$ $3.6001$ $41.3945$ $0.9693$ $1.7049$ $1.5391$ $36.7449$ $0.9367$ $5.1862$ $5.2610$
    FusionNet (Deng et al., 2020) $36.4478$ $0.9567$ $4.4273$ $3.6040$ $41.0195$ $0.9673$ $1.7542$ $1.5953$ $36.2621$ $0.9311$ $5.4256$ $5.5555$
SatFusion     Sharpen     MISR
PNN SRCNN $36.4483$ $0.9580$ $4.3532$ $3.6277$ $41.0877$ $0.9685$ $1.7861$ $1.5841$ $35.7310$ $0.9247$ $5.2484$ $5.9422$
HighRes-Net $36.0293$ $0.9518$ $4.8236$ $3.8285$ $41.6078$ $0.9692$ $1.6793$ $1.5126$ $35.9192$ $0.9256$ $5.2069$ $5.8500$
RAMS $37.0329$ $0.9608$ $4.1276$ $3.4447$ $41.7379$ $0.9693$ $1.6513$ $1.4934$ $36.2884$ $0.9291$ $5.1132$ $5.5815$
TR-MISR $36.7192$ $0.9591$ $4.3381$ $3.5696$ $41.7699$ $0.9695$ $1.6360$ $1.4894$ $36.0572$ $0.9275$ $5.1791$ $5.7769$
DiCNN SRCNN $37.6085$ $0.9654$ $3.8363$ $3.2232$ $42.3337$ $0.9731$ $1.4309$ $1.4096$ $36.8326$ $0.9397$ $5.0185$ $5.2307$
HighRes-Net $37.8380$ $0.9660$ $3.7481$ $3.1560$ $42.6423$ $0.9740$ $1.3664$ $1.3712$ $37.7222$ $0.9480$ $4.5623$ $4.6345$
RAMS $37.9592$ $0.9670$ $3.7040$ $3.1087$ $41.9437$ $0.9694$ $1.4461$ $1.4873$ $38.0532$ $0.9515$ $4.5044$ 4.4589
TR-MISR $37.8946$ $0.9669$ $3.7303$ $3.1411$ $43.0302$ $0.9755$ $1.3201$ $1.2979$ $37.2068$ $0.9450$ $4.6157$ $4.9595$
MSDCNN SRCNN $35.7101$ $0.9476$ $4.9176$ $3.9733$ $40.7670$ $0.9661$ $1.8893$ $1.6714$ $35.4558$ $0.9207$ $5.5356$ $6.1120$
HighRes-Net $35.8939$ $0.9463$ $5.0362$ $3.9184$ $40.8732$ $0.9668$ $1.8991$ $1.6453$ $35.6254$ $0.9226$ $5.2722$ $6.0208$
RAMS $36.2144$ $0.9494$ $4.7771$ $3.8197$ $41.6595$ $0.9689$ $1.6937$ $1.5094$ $35.7382$ $0.9239$ $5.2642$ $5.9819$
TR-MISR $36.2069$ $0.9522$ $4.7088$ $3.7671$ $41.2276$ $0.9657$ $1.7653$ $1.5762$ $35.6936$ $0.9242$ $5.2186$ $5.9742$
DRPNN SRCNN $36.8788$ $0.9595$ $4.2331$ $3.5196$ $42.4993$ $0.9738$ $1.4364$ $1.3783$ $37.4546$ $0.9452$ $4.8336$ $4.8230$
HighRes-Net $37.2064$ $0.9619$ $4.0026$ $3.3926$ $43.1166$ $0.9760$ $1.3685$ $1.2851$ $37.8682$ $0.9500$ $4.6161$ $4.5614$
RAMS $37.2588$ $0.9618$ $4.0436$ $3.3597$ $43.5475$ $0.9775$ $1.3406$ $1.2095$ $37.9408$ $0.9500$ $4.5958$ $4.5190$
TR-MISR $37.0312$ $0.9603$ $4.1118$ $3.4233$ $43.5426$ $0.9778$ $1.3176$ $1.2127$ $37.9545$ $0.9506$ $4.5695$ $4.4999$
FusionNet SRCNN $37.6541$ $0.9656$ $3.7718$ $3.2289$ $42.7206$ $0.9749$ $1.3651$ $1.3452$ $37.7903$ $0.9496$ $4.6609$ $4.6268$
HighRes-Net $38.1041$ $0.9674$ $3.6253$ $3.0475$ $43.3822$ $0.9777$ $1.3050$ $1.2506$ $37.9892$ $0.9520$ 4.4420 $4.5124$
RAMS 38.3352 0.9698 3.5193 2.9545 $43.4439$ $0.9774$ $1.2982$ $1.2333$ 38.0487 0.9526 $4.4452$ $4.4753$
TR-MISR $38.1832$ $0.9682$ $3.5763$ $3.0154$ 43.8911 0.9793 1.2683 1.1717 $37.8690$ $0.9514$ $4.4463$ $4.5625$
  • The bold entities represent the best results in each sub-block.
Figure 4. (left) Workflow of the conventional Wald protocol for dataset construction; (right) Workflow of the Wald protocol with introduced noise for dataset construction. :
Figure 5. Quantitative comparison of PSNR for the fused images on the WorldStrat real-world dataset. :
Figure 6. Qualitative comparison of error maps between the fused and ground-truth images. :
5.1. Experimental Setup
Dataset. We conduct experiments on the WorldStrat (Cornebise et al., 2022), WV3, QB, and GF2 datasets (Deng et al., 2022). WorldStrat data comes from real satellites without removing low-quality images, thus better reflecting real-world scenarios. Each location in this dataset consists of multiple multi-temporal LRMS images, one HRPAN image, and one HRMS image, which are commonly used for MISR network training. For the WV3, QB, and GF2 datasets, each location includes one HRPAN image and one LRMS image. Previous works commonly employed these datasets to construct synthetic datasets via the Wald protocol (Wald et al., 1997) for pansharpening network training.
We conduct experiments on both the real dataset and synthetic datasets constructed using the Wald protocol. It is worth noting that SatFusion takes multiple multi-temporal LRMS images and a single HRPAN image as input. However, as shown in Figure 4(left), prior pansharpening works use Wald protocol to construct synthetic lrmslrms and PanPan images as input, which has inherent idealized limitations (Section 2.2). As shown in Figure 4(right), we introduce pixel shifts and noise perturbations into the Wald protocol to construct synthetic multi-temporal images $\{lrms_{t}\}_{t=1}^{T}$, which are then combined with the PanPan image as the input of SatFusion. Experiments are conducted on both real datasets and synthetic datasets with added noise perturbations, making them more consistent with the data conditions in Sat-IoT scenarios.
Evaluation Methods and Metrics. We compare SatFusion against several MISR networks and pansharpening networks. We adopt PSNR, SSIM, SAM (Yuhas et al., 1992), and ERGAS (Alparone et al., 2007) as quantitative evaluation metrics, and additionally provide reconstructed images and error maps for qualitative evaluation.
Implementation Details. We ensure consistency in the experimental setup, meaning that when comparing SatFusion with baseline methods, training parameters (e.g., epochs, batch size, optimizer, etc.) and data parameters (e.g., $H$, $W$, $C_{ms}$, $\gamma$, $T$, etc.) are kept the same. For more details on the experimental parameter settings, please refer to Appendix E.
5.2. Main Experiments
We conduct experiments on both real and synthetic datasets to answer RQ1 and RQ2.
On the real dataset, we validate that the SatFusion framework is beneficial for enhancing both MISR and Pansharpening networks. We conduct experiments on the original WorldStrat dataset, where SatFusion takes $\{lrms_{t}\}_{t=1}^{T=8}$ and PanPan as inputs,   the MISR network takes $\{lrms_{t}\}_{t=1}^{T=8}$ as input, and the Pansharpening network takes a randomly selected single lrmslrms and PanPan as inputs.
As shown in Table 1(a), SatFusion achieves notable improvements over both MISR and Pansharpening networks in quantitative metrics. Specifically, compared with MISR, the average PSNR is improved by 27.1% and ERGAS by 73.6%; compared with Pansharpening, the average PSNR and ERGAS are improved by 2.6% and 13.9%, respectively. On the WorldStrat test set, SatFusion demonstrates a clear enhancement in the reconstructed image quality, indicating its capability to adaptively leverage the strengths of both MISR and Pansharpening methods. We visualize the fused results and their corresponding PSNR values in Figure 5.
On the synthetic dataset, we validate that the SatFusion framework is beneficial for enhancing Pansharpening networks. Using the WorldStrat, WV3, QB, and GF2 datasets, we construct synthetic multi-temporal images $\{lrms_{t}\}_{t=1}^{T}$ with pixel shifts and noise disturbances introduced through the Wald protocol. SatFusion takes $\{lrms_{t}\}_{t=1}^{T=8}$ and PanPan as inputs, while the Pansharpening network takes a randomly selected lrmslrms and PanPan as inputs.
As shown in Table 2 and Table 1(b), SatFusion significantly improves the quality of the fused images on the test sets of the synthetic datasets. We visualize the fused images and the error maps between the SRSR and the GTGT in Figure 6, verifying its ability to adaptively fuse multi-source features in scenarios with complex disturbances. This further demonstrates the effectiveness and advantages of SatFusion in integrating MISR and Pansharpening networks.
6. Discussion
6.1. Practical Significance of SatFusion
Enhancing image quality. As an end-to-end unified framework, SatFusion unifies the advantages of existing fusion paradigms to enhance reconstruction quality. Unlike MISR, which relies solely on multi-temporal data, it leverages high-resolution spatial details from the HRPAN image. Unlike conventional pansharpening that interpolates a single LRMS image, SatFusion achieves alignment by incorporating HR spatial features generated from complementary multi-temporal observations. This design ensures robustness under motion blur and large modality gaps, effectively bridging the gap between redundant low-quality data and the demand for high-quality imagery in real-world Sat-IoT applications.
Reducing image redundancy. Traditional approaches typically store and transmit multiple images across temporal phases, leading to redundant transmission of overlapping information. In contrast, SatFusion consolidates multiple low-quality images into a single high-quality output, preserving information integrity while avoiding redundant storage and transmission. This naturally leads to a degree of data compression.
Assuming each pixel in every channel occupies one unit of storage space, let NN denote the number of locations. Then the input data size is $D_{\text{input}}=N\times\Big(T\cdot\tfrac{H}{\gamma}\cdot\tfrac{W}{\gamma}\cdot C_{ms}\;+\;H\cdot W\cdot 1\Big),$ and the output data size is $D_{\text{output}}=N\times(H\cdot W\cdot C_{ms}).$ While redundancy reduction is guaranteed, the total physical data volume depends on specific conditions. Because the fused output typically has higher spatial resolution and more channels, the enhanced information richness may increase total data size. Only when $T$ is sufficiently large does the compression benefit outweigh this increase, satisfying $D_{\text{output}}<D_{\text{input}}$. As shown in Figure 7, this advantage becomes increasingly pronounced in Sat-IoT scenarios characterized by abundant multi-temporal data.
Figure 7. Difference in data volume between input and output images ($D_{\text{in}}-D_{\text{out}},H(W)=1024,\gamma=4$) :
6.2. Future Prospects
SatFusion takes $\{lrms_{t}\}_{t=1}^{T}$ and Pan{Pan} as inputs, where the multi-temporal $\{lrms_{t}\}_{t=1}^{T}$ images not only provide complementary information but also contribute to learning deep features aligned with PanPan. However, in real-world scenarios, multi-temporal PAN images also exist but were not utilized in our design. Intuitively, incorporating multi-temporal PAN images through novel architectural designs could further improve fusion performance.
Moreover, SatFusion constitutes the first unified framework in the Sat-IoT domain that jointly optimizes high-quality image generation from both multi-temporal and multi-source data. Future research could extend this direction by exploring more sophisticated fusion architectures, adaptive loss balancing mechanisms, and lightweight model, further bridging the gap between theoretical fusion methods and real-world satellite applications.
7. Conclusion
In this paper, we propose SatFusion, the first end-to-end unified framework that leverages multi-temporal and multi-source images for collaborative optimization to generate high-quality images, overcoming the limitations of task separation in existing methods. Quantitative and qualitative experiments conducted on the WorldStrat, WV3, GF2, and QB benchmark datasets demonstrate that SatFusion significantly improves fusion quality. Moreover, SatFusion maintains stable performance in challenging scenarios such as image blurring and large modality discrepancies. By alleviating the contradiction between redundant low-quality data and the demand for high-quality data, SatFusion better aligns with real-world Sat-IoT scenarios and holds strong potential for practical applications.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
Appendix A Hyperparameter Experiments
We address RQ3 by varying two key hyperparameters.
We first change the number of input multi-temporal images $T$ on the original WorldStrat dataset. As shown in Figure 8, the quality of image fusion improves as the number of multi-temporal images increases. In particular, the evaluation metrics show significant improvement when the number of images increases in the early stage, indicating that incorporating more multi-temporal data is beneficial for the model performance. However, when the number of images continues to increase, the metrics tend to stabilize. We attribute this to two reasons: (1) limitations of the model architecture design, and (2) the fact that introducing more multi-temporal data brings not only complementary information but also more noise (An et al., 2022), which constrains further performance gains.
Figure 8. Changes in metrics for different fusion methods as the number of input multi-temporal images $T$ increases. :
We also vary the super-resolution scaling factor $\gamma$ on the original WorldStrat dataset. As shown in Table 3, SatFusion achieves the best performance at scaling factors $\gamma=2,5$ (In Table  1, $\gamma\approx 3$). Moreover, for relatively small scaling factors, the fused super-resolved MS images SRSR produced by SatFusion exhibit more stable quality.
Table 3. Quantitative results at scales $\uparrow\times 2$ and $\uparrow\times 5$.
[…]
MISR HighRes-Net 37.905 9437.905\,94 0.888 560.888\,56 2.094 592.094\,59 8.316 368.316\,36
RAMS 38.163 9938.163\,99 0.892 450.892\,45 2.082 822.082\,82 7.909 037.909\,03
Sharpen PNN 47.299 6647.299\,66 0.987 790.987\,79 2.059 952.059\,95 2.459 642.459\,64
Pan-Mamba 45.786 2745.786\,27 0.986 050.986\,05 2.994 142.994\,14 2.994 412.994\,41
SatFusion HighRes-Net+PNN 48.1212 0.9904 1.8379 2.1995
RAMS+Pan-Mamba 47.577 7847.577\,78 0.987 760.987\,76 1.936 951.936\,95 2.485 602.485\,60
[…]
MISR SRCNN 35.518 9035.518\,90 0.854 300.854\,30 2.556 422.556\,42 8.594 368.594\,36
TR-MISR 35.668 6435.668\,64 0.856 770.856\,77 2.621 942.621\,94 8.385 408.385\,40
Sharpen PanNet 44.712 6244.712\,62 0.980 570.980\,57 2.440 562.440\,56 3.079 373.079\,37
INNformer 45.514 4845.514\,48 0.984 290.984\,29 2.493 362.493\,36 2.763 182.763\,18
SatFusion SRCNN+PanNet 46.158 6646.158\,66 0.9888 2.180 092.180\,09 2.540 532.540\,53
TRMISR+INNformer 46.3109 0.988 700.988\,70 2.0627 2.4422
  • The bold entities represent the best results.
Appendix B Ablation Studies
We address RQ4 by conducting ablation experiments on both the loss function components and the Fusion Composition module.
We design different weighted combinations of loss functions to supervise the reconstruction results, jointly optimizing both texture and spectral features of the images. We perform ablation studies on the loss components using the original WorldStrat dataset. As shown in Table 4, the worst metrics are highlighted in bold. Removing any single loss term leads to a significant degradation in certain evaluation metrics, confirming that our multi-loss optimization effectively balances performance and perceptual quality.
Table 4. Quantitative results with different loss functions.
[…]
SRCNN +PNN 47.231 8547.231\,85 0.989 780.989\,78 2.008 162.008\,16 2.371 852.371\,85
47.181 6847.181\,68 0.9896 2.1779 2.411 682.411\,68
46.8231 0.990 380.990\,38 1.919 291.919\,29 2.5160
47.308 2147.308\,21 0.990 790.990\,79 2.086 912.086\,91 2.387 432.387\,43
HighRes-Net +PanNet 47.391 1047.391\,10 0.990 220.990\,22 1.950 811.950\,81 2.323 232.323\,23
47.565 7047.565\,70 0.990 340.990\,34 1.9903 2.259 232.259\,23
46.7710 0.988 930.988\,93 1.880 971.880\,97 2.5688
47.427 4147.427\,41 0.9875 1.986 631.986\,63 2.287 592.287\,59
RAMS +INNformer 47.464 9547.464\,95 0.990 210.990\,21 1.887 711.887\,71 2.363 472.363\,47
47.553 5247.553\,52 0.987 120.987\,12 2.098 932.098\,93 2.378 222.378\,22
47.0930 0.986 860.986\,86 1.897 411.897\,41 2.439 342.439\,34
47.629 2747.629\,27 0.9862 2.3441 2.5325
TR-MISR +Pan-Mamba 47.093 3347.093\,33 0.987 930.987\,93 2.005 482.005\,48 2.487 952.487\,95
47.308 6047.308\,60 0.986 000.986\,$0$ 2.209 882.209\,88 2.509 512.509\,51
46.4710 0.9832 2.003 972.003\,97 2.7849
47.231 1047.231\,10 0.985 440.985\,44 2.4206 2.756 362.756\,36
  • The bold entities represent the worst results.
The Fusion module consists of residual connections and convolutional adjustment operations. Compared with synthetic datasets, the real datasets exhibit more noticeable spectral differences between the multi-temporal images and the HRMS images. We conduct ablation on the convolutional adjustment operation using the original WorldStrat dataset. As shown in Table 5, removing the convolutional adjustment operation results in a drop in most metrics and visible color distortions, thereby reducing perceptual quality. We also observe that for a few methods (e.g., Pan-Mamba), the convolutional adjustment operation has limited influence. However, as a general framework that should adapt to a wide range of prior works, including the convolutional adjustment operation is necessary.
Table 5. Ablation study on the effect of convolutional adjustment operation.
[…]
HighRes-Net+PNN 47.3835 0.9914 1.9368 2.3304
47.199 3847.199\,38 0.989 070.989\,07 1.971 371.971\,37 2.343 262.343\,26
HighRes-Net+INNformer 47.4021 0.9902 1.8958 2.3094
47.328 0447.328\,04 0.989 730.989\,73 1.941 301.941\,30 2.320 982.320\,98
TR-MISR+INNformer 47.1728 0.9863 1.8978 2.4066
46.798 7646.798\,76 0.985 080.985\,08 2.001 452.001\,45 2.511 962.511\,96
RAMS+INNformer 47.4650 0.9902 1.8877 2.3635
47.138 5047.138\,50 0.987 590.987\,59 1.941 101.941\,10 2.403 202.403\,20
TR-MISR+PanNet 47.1593 0.9901 1.900 061.900\,06 2.3601
47.129 6047.129\,60 0.989 750.989\,75 1.8937 2.365 752.365\,75
HighRes-Net+Pan-Mamba 46.311 1546.311\,15 0.988 670.988\,67 2.0229 2.6026
46.3500 0.9909 2.239 292.239\,29 2.668 732.668\,73
TR-MISR+Pan-Mamba 47.0933 0.987 930.987\,93 2.0055 2.487 952.487\,95
47.076 7547.076\,75 0.9899 2.051 402.051\,40 2.4806
  • The bold entities represent the better results.
Appendix C Robustness Experiments
We address RQ5 by varying the noise level to evaluate model robustness.
To further validate the robustness and generalization ability of the proposed method in complex remote sensing scenarios, we conducted experiments on a synthetic dataset by introducing a control parameter $\varepsilon$ to jointly adjust the intensity of noise perturbations, thereby simulating the instabilities commonly observed in real multi-temporal remote sensing imagery.
[…]
As shown in Figure 9, the experimental results indicate that as $\varepsilon$ increases, the degradation of fused image quality in SatFusion is more gradual, suggesting that the method can maintain stable performance even under challenging conditions such as severe image blurring and significant cross-modal discrepancies.
Figure 9. PSNR comparison under different perturbation levels $\varepsilon$. SatFusion incorporating multi-temporal information consistently show higher reconstruction quality, demonstrating stronger robustness. :
Appendix D Details of Loss Computation
As emphasized in the main text, we ensured consistency in loss computation throughout the experiments so that the optimization of evaluation metrics directly demonstrates the effectiveness of the proposed architecture. This section provides a detailed description of the loss computation process.
We conducted experiments on the WorldStrat, WV3, QB, and GF2 datasets. Previous works commonly use the WV3, QB, and GF2 datasets to construct artificial datasets via the Wald protocol for training Pansharpening networks. In our experiments validating that integrating Pansharpening networks into SatFusion can improve fusion quality and robustness, we modified the DLPan-Toolbox codebase (Deng et al., 2022). This codebase, which evaluates the effectiveness of Pansharpening network structures using WV3, QB, and GF2, integrates networks such as FusionNet and DRPNN. Based on this codebase, we built the SatFusion framework by integrating MISR networks with Pansharpening networks, while also adjusting the input data (as shown in Fig. 6(b)). In these experiments, SatFusion and Pansharpening networks adopted exactly the same experimental configurations (training epochs, batch size, optimizer, etc.) and the same loss functions as the original Pansharpening network training, to demonstrate the generalizability of SatFusion’s ability to enhance prior work.
Previous studies often employ the WorldStrat real dataset for MISR network training. In experiments on this dataset validating the integration of Pansharpening networks into SatFusion, we followed prior work by performing image adjustment before loss computation to address global brightness bias (An et al., 2022; Deudon et al., 2020; Salvetti et al., 2020). Specifically, we corrected the overall brightness of the SatFusion output HRMS image to match that of the ground truth HRMS image by computing the mean bias channel-wise:
$\mu_{GT}=\frac{1}{H\times W}\sum_{i=1}^{H}\sum_{j=1}^{W}GT_{i,j},\quad\mu_{SR}=\frac{1}{H\times W}\sum_{i=1}^{H}\sum_{j=1}^{W}SR_{i,j},$
$SR=SR+\mu_{GT}-\mu_{SR}.$
Here, $\mu_{GT}$ denotes the overall brightness of the ground truth HRMS image, while $\mu_{SR}$ denotes the overall brightness of the SatFusion output HRMS image. The brightness compensation is applied channel-wise.
To handle potential image misalignment, SatFusion crops both GTGT and SRSR to compensate for spatial shifts and returns the minimum loss value (Cornebise et al., 2022). Cropping is implemented using a Lanczos convolution kernel. Lanczos is a high-quality image resampling filter based on the sinc function. It achieves a balance between computational efficiency and anti-aliasing by truncating the infinite support of the sinc function with a window function, thereby reducing ringing artifacts while preserving sharpness. The initial offset $x$ is generated from $p_{x}$ (absolute shift), stepstep (shift stride), and $a$ (kernel lobes):
$r=start,start+1,\dots,end,$
$d=-px,-px+step,-px+2step,\dots,px,$
$x_{i}=r_{1}-d_{i},r_{2}-d_{i},\dots,r_{width}-d_{i},$
where $start=-a-p_{x}$, $end=a+p_{x}$, and widthwidth is the length of $r$. A sine filter is then applied to generate a 1D kernel $L$, which is extended to a 2D Lanczos kernel for image shifting. For each shifted ground truth image $GT_{i}^{\prime}$, we align it with the cropped $SR^{\prime}$ and compute the loss, selecting the minimum as the final result:
$L_{i}=a\cdot\mathrm{sinc}(\pi\cdot x_{i})\cdot\mathrm{sinc}\left(\frac{\pi\cdot x_{i}}{a}\right),\quad Lanczos=L\otimes L,$
$GT_{i}^{\prime}=Lanczos_{i}(GT),\quad SR^{\prime}=patch(SR),$
$loss=\min\,Loss(GT_{i}^{\prime},SR^{\prime}).$
Here, $patch(\cdot)$ is a cropping function that adjusts the image size to match $GT_{i}^{\prime}$. LossLoss refers to the composite weighted loss introduced in Section 4.3, with weights $\lambda_{1}=0.3$, $\lambda_{2}=0.3$, $\lambda_{3}=0.2$, and $\lambda_{4}=0.2$.
Appendix E Details of Experimental Parameter Settings
As emphasized in the main text, we ensured the consistency of experimental configurations, i.e., SatFusion and baseline methods were compared under the same training and data parameter settings. This section provides the detailed parameter configurations used in the experiments.
The WorldStrat dataset provides MISR network codes as baselines, including SRCNN and HighRes-Net. In addition, we incorporated other representative MISR methods (RAMS, TR-MISR) and representative Pansharpening methods (PNN, PanNet, INNformer, Pan-Mamba). For experiments conducted on this dataset, the original parameters in the source codebase were preserved (Cornebise et al., 2022). Specifically, the training parameters were set as: batch size = 16, optimizer = Adam, and maximum epochs = 20. The data parameters included consistent train/validation/test splits, with $H=W=156,\tfrac{H}{\gamma}=\tfrac{W}{\gamma}=50\ (\gamma\approx 3),C_{ms}=3,T=8,C_{\text{hidden\_encode}}=128,C_{MISR}=128$. Here, to ensure that $C_{MISR}$ is divisible by ${\gamma}^{2}$, the parameter $\gamma$ of the PixelShuffle Block is set to 2. In all other cases, the PixelShuffle Block parameter $\gamma$ is identical to the image spatial upscaling factor $\gamma$.
Extensive studies of Pansharpening networks have been conducted on the WV3, QB, and GF2 datasets. The DLPan-Toolbox codebase provides corresponding implementations of Pansharpening methods. We used five representative baselines: PNN, DiCNN, MSDCNN, DRPNN, and FusionNet, and incorporated MISR networks into SatFusion. In experiments on these datasets, the data parameters were consistent with the source codebase, while training parameters were set individually for different Pansharpening methods (Deng et al., 2022). Taking WV3 as an example, the experimental configurations (with reduced epochs compared to the original codebase for efficiency, in fact the model has already converged) were as follows. For training and validation: $H=W=64,\tfrac{H}{\gamma}=\tfrac{W}{\gamma}=16\ (\gamma=4),C_{ms}=8,T=8,C_{\text{hidden\_encode}}=128,C_{MISR}=128.$ For testing: $H=W=256,\tfrac{H}{\gamma}=\tfrac{W}{\gamma}=64\ (\gamma=4),$ with other parameters consistent with those of the training and validation sets.
Table 6. Training parameters for different methods.
Parameters PNN DiCNN MSDCNN DRPNN FusionNet
Epochs 1000 1000 500 500 400
Batch Size 64 64 64 32 32
Optimizer SGD Adam Adam Adam Adam
Loss Function $l_{2}$ $l_{2}$ $l_{2}$ $l_{2}$ $l_{2}$
All experimental code has been open-sourced. Experiments were conducted on a server equipped with eight NVIDIA RTX 4090 GPUs to ensure reproducibility.